Sol Summary Report – February 16th

Crew 275 Sol Summary Report 16Feb2023

Sol: 4

Summary Title: “HabCom, do you copy?”

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today was certainly the most intense day since the beginning of the simulation. We woke up a bit earlier and shortened our daily sport session, so that we could get ready for a relatively technical EVA. With temperatures getting as low as -17°C/1°F, the three EVA crewmembers had to wear very warm clothes that still allowed them to move freely so that they could proceed with the deployment of the MegaAres instrument. After a three-hour EVA, the first step of the installation has now been accomplished in an area free of snow less than 250 meters away from the Hab. The next step will be to connect the instrument to an external battery during another EVA and install a meteorological mast for contextualization of the data. We spent a large part of the afternoon rehearsing for the deployment of this mast. The cold weather also affected the power supply systems, with a short blackout in the morning, which was fortunately soon fixed by mission support. In the evening we continued to be very careful with our power consumption to be gentle with the generator. On the bright side, we are learning from our mistakes, getting better and better with the bread-making machine, which is fundamental for French people.

Look Ahead Plan:

Should the weather and the ground conditions allow it, we would like to continue the deployment process of the instruments dedicated to atmospheric measurements. To do this, we are planning an EVA tomorrow morning to bring power supply to MegaAres, test it, and install the meteorological mast right next to it. Due to the weather forecast, the EVA dedicated to the deployment of LOAC may be further delayed.

On Saturday, we would like to make a first EVA dedicated to photogrammetry, using drones.

In the meantime, we will proceed with the deployment of our environmental sensors throughout the campus.

Anomalies in work: Issues with the generator, leading to a blackout in the morning and the necessity to limit power consumption more than usual during the rest of the day. In addition, we observed that the robotic observatory has been left in open position since at least this morning.

Weather: Sunny, low -17°C/1°F, high 1°C/34°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations

Report, Journalist Report, HSO Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site